A 6.82 kg bowling ball moves at 4.29 m/s. How fast must a 2.36 g Ping-Pong ball move so that the two balls have the same kinetic
energy?
Answer in units of m/s
1
Expert's answer
2012-11-29T08:00:38-0500
The kinetic energy is K=(mV^2)/2
Let x m/s is the velocity of Ping- Pong ball. If the kinetic energies are the
same then:
6.82*(4.29)^2/2=0.00236*x^2/2
x^2=6.82*(4.29)^2/0.00236=53184.7
x=230.6 m/s